OpenStudy (anonymous):

Solve for y. 7y - 6y - 10 = 13 y = -23 y = -3 y = 3 y = 23

OpenStudy (anonymous):

y=23 is there any need of an explanation?

OpenStudy (anonymous):

y=23

OpenStudy (anonymous):

\(\sf 7y - 6y - 10 = 13\\7y-6y=13+10\\y=23\) 2nd step: collect like terms 3rd step: simplify